OpenStudy (anonymous):

If cos2 x + 2cosx - 2 = 0, then cosx = _____. -1 ±√3 1 ±√3 ±√3 1

OpenStudy (solomonzelman):

Can you solve? \(\large\color{black}{ \displaystyle k^2+2k-2=0 }\)

OpenStudy (solomonzelman):

You can do this: \(\large\color{black}{ \displaystyle k^2+2k-2=0 }\) \(\large\color{black}{ \displaystyle k^2+2k-2\color{red}{+3}=0 \color{red}{+3} }\) \(\large\color{black}{ \displaystyle k^2+2k+1=3 }\) \(\large\color{black}{ \displaystyle \left(k+1\right)^2=3 }\)

OpenStudy (solomonzelman):

via «completing the square» method.